# Locally recoverable algebro-geometric codes from projective bundles

### v2

@@ Abstract.
A code is locally recoverable when each symbol in one of its code words can be reconstructed as a function of $r$ other symbols. We use bundles of projective spaces over a line to construct locally recoverable codes with availability; that is, evaluation codes where each code word symbol can be reconstructed from several disjoint sets of other symbols. The simplest case, where the code’s underlying variety is a plane, exhibits noteworthy properties: When $r=1$, $2$, $3$, they are optimal; when $r\geq 4$, they are optimal with probability approaching $1$ as the alphabet size grows. Additionally, their information rate is close to the theoretical limit. In higher dimensions, our codes form a family of asymptotically good codes.
@@ 1.2. Results
Let $r=1$, $2$, or $3$ and let $b$ satisfy $3\leq b\leq\frac{q}{r+1}$. The code $\mathcal{C}$ is optimal and locally recoverable with locality $r$. Its parameters $[n,k,d]_{q}$ are
@@
Let ${\mathbb{F}}_{q}$ be a finite field of cardinality $q$. Fix positive integers $\alpha$, $\beta$, $b$, and $t$ such that $b\geq\alpha+1$. There exists an integer $m_{0}=m_{0}(\alpha,\beta,b,t,q)$ such that, for each $m\geq m_{0}$, there is a locally recoverable code over ${\mathbb{F}}_{q}$ with locality $r=\binom{\beta+m}{\beta}$ and availability $t$, and parameters $[n,k,d]_{q}$ satisfying
@@
The cardinality of a set $S$ will be denoted by #$\#S$. A code will be denoted by a calligraphic letter ${\mathcal{C}}$, and its dual code will be denoted by ${\mathcal{C}}^{\bot}$. All the codes we handle are linear evaluation codes ${\mathcal{C}}\coloneqq\mathrm{im}\operatorname{ev}_{{\mathcal{P}}}$ with ${\mathcal{P}}$ a finite set of points. A code word of ${\mathcal{C}}$ will be denoted by a bold lowercase letter $\mathbf{c}$. Scalars in ${\mathbb{F}}_{q}$ will be denoted by a lowercase $x$, and elements of ${\mathbb{F}}_{q}^{n}$ for $n>1$ will be denoted by a bold lowercase $\mathbf{x}$. The $n$-dimensional affine space over ${\mathbb{F}}_{q}$ is denoted by ${\mathbb{A}}^{n}$, and its Cartesian coordinates are denoted by ${\mathbb{A}}_{x_{1},\dots,x_{n}}^{n}$. The $n$-dimensional projective space over ${\mathbb{F}}_{q}$ is denoted by ${\mathbb{P}}^{n}$, its homogeneous coordinates are denoted by ${\mathbb{P}}_{\mathbf{x}}^{n}$ where $\mathbf{x}=[x_{0},\dots,x_{n}]$. When we write ${\mathbb{P}}_{x_{0},\dots,x_{n-1}}^{n}$, the coordinates $x_{0},\dots,x_{n-1}$ are the affine coordinates in the chart $D_{+}(x_{n})$ where the last homogeneous coordinate $x_{n}$ is nonzero. Given a projective variety $X$ over ${\mathbb{F}}_{q}$, we denote by $X({\mathbb{F}}_{q})$ the ${\mathbb{F}}_{q}$-rational points of $X$.
@@ 2.1. Linear locally recoverable codes
A code ${\mathcal{C}}$ is said to be locally recoverable (LRC) with locality $r$ if for each symbol $c_{i}$ in a code word $\mathbf{c}=(c_{1},\dots,c_{n})\in{\mathcal{C}}$, there exists a recovery set $R_{i}\subset\{1,\dots,n\}\setminus\{i\}$ with #$\#R_{i}\leq r$ such that $c_{i}$ is a function of the symbols $\{c_{j}\}_{j\in R_{i}}$. In particular, if the $i$-th coordinate of $\mathbf{c}$ is lost, it can be recovered by accessing $\leq r$ other coordinates in $\mathbf{c}$. Trivially, every linear code is a LRC with locality $r=k$. Our convention of locality throughout this paper is also referred to as all-symbol locality in the literature [BargTamoVluadact17]. An LRC ${\mathcal{C}}$ with locality $r$ is said to have availability $t$ if for every $c_{i}$ in a code word $\mathbf{c}=(c_{1},\dots,c_{n})\in{\mathcal{C}}$ there exist $t$ disjoint recovery sets $R_{i,\ell}\subset\{1,\dots,n\}\setminus\{i\}$ with #$\#R_{i,\ell}\leq r$ such that $c_{i}$ is a function of the symbols $\{c_{j}\}_{j\in R_{i,\ell}}$ for $1\leq\ell\leq t$.
The parameters of an LRC ${\mathcal{C}}$ are denoted $[n,k,d;r,t]_{q}$, or simply $[n,k,d;r]_{q}$ when $t=1$. They are constrained by relations like the Singleton-type bound for the minimum distance
@@
An LRC ${\mathcal{C}}$ with parameters $[n,k,d;r]_{q}$ whose minimum distance achieves equality in (2.1) is said to be optimal. When an LRC ${\mathcal{C}}$ with parameters $[n,k,d;r]_{q}$ has the property that each code word is partitioned into sets of $r+1$ elements where recoverability takes place within each set, as is often the case with algebro-geometric locally recoverable codes, a particularly simple proof of (2.1) was given in [SalgadoVarillyAlvaradoVoloch21]*Theorem I.3.
Let $\mathscr{C}=\{{\mathcal{C}}_{i}\}_{i=1}^{\infty}$ be a family of codes (not necessarily LRCs) and denote by $[n_{i},k_{i},d_{i}]_{q}$ the parameters of ${\mathcal{C}}_{i}$ for all $i\geq 1$. We say that $\mathscr{C}$ is a family of asymptotically good codes when
@@ 2.2. Algebro-geometric evaluation codes
Algebraic Geometry supplies an abundance of constructions for linear codes under the framework of evaluation codes. To specify such a code one needs a triple $(X,{\mathcal{P}},V)$, where $X$ is a quasi-projective variety over a finite field ${\mathbb{F}}_{q}$, ${\mathcal{P}}$ is a set of $n$ points in $X({\mathbb{F}}_{q})$, and $V$ is a finite-dimensional subspace of the function field of $X$. From this data, we construct a linear code as the image of the evaluation map
${\operatorname{ev}_{\mathcal{P}}\colon}$${V}$${{\mathbb{F}}_{q}^{n},}$${f}$${(f(P_{1}),\dots,f(P_{n})).}$
@@ 3.2. Relations to existing codes
for a nonnegative integer $N$. Our codes consider mostly the case where $N=b-2$, though we look at smaller values of $N$ in Section 3.7. The key difference between their codes and ours is the choice of points ${\mathcal{P}}$ for code evaluation. We all consider a set of points ${\mathcal{P}}$ partitioned into $b$ batches $A_{1},\dots,A_{b}$ consisting of $r+1$ distinct points. However, in [SalgadoVarillyAlvaradoVoloch21]*Section III.B each batch is additionally constrained to satisfy an extra algebraic relation: If $A_{i}=\{(x_{i},y_{i,j})\}_{j=1}^{r+1}$, then there exists $g(x)\in{\mathbb{F}}_{q}[x]$ a polynomial of degree $r+1$ such that $y_{i,j}=g(x_{i})$ for all $1\leq i\leq b$ and $11\leq j\leq r+$ 1. Tamo and Barg’s construction in [TamoBarg14] uses the polynomial $g(x)=x^{r+1}$, and more generally discusses the concept of good polynomials $g(x)$. The polynomial $g(x)$ affords better control of the minimum distance of the resulting codes. Without this kind of control, we can only prove that the codes constructed in Section 3.1 are optimal when $b\geq 3$ and $r=1$, $2$ or $3$; see Theorem 3.13. However, when $r\geq 4$, our codes in Section 3.1 are provably optimal with probability approaching $1$ for uniformly random choices of points ${\mathcal{P}}$ as $q\to\infty$; see Theorem 3.15. Although our codes for $r\geq 4$ are not always optimal (see Section 4), our investigation suggests that the use of good polynomials, like $g(x)=x^{r+1}$, imposes a serious constraint on the universe’s supply of optimal LRCs.
Our codes also share superficial similarities with those of Munuera and Tenório [MunueraTenorio18]*Section 2.2, which is not surprising, as their codes generalize those in [TamoBarg14, BargTamoVluadact17]. However, our codes are neither a special case of the Munuera–Tenório construction, nor is there a clear common refinement of both constructions. To mimic our codes in the notation of [MunueraTenorio18], one would need to take $m=2$, $t=1$, $\phi_{1}=x_{1}\eqqcolon x$, $\phi_{2}=x_{2}\eqqcolon y$, #$\#{\mathcal{S}}=b$, and $V_{i}=\{a(x)\in{\mathbb{F}}_{q}[x]:\deg a(x)\leq b-2\}$ for all $10\leq i\leq r-$ 1. This would force $r=q-1$ in their set-up (a restriction we do not impose), and the set of evaluation points ${\mathcal{P}}\subset{\mathbb{A}}^{2}_{x,y}({\mathbb{F}}_{q})$ would contain $b(r+1)=bq$ points in $b$ batches with $r+1=q$ overlapping $y$ coordinates. In our construction, it is essential that all the $b(r+1)$ points have distinct $y$ coordinates; see Remarks 3.1(3) and Figure 1.
@@ 3.3. Dimension and locality of ${\mathcal{C}}$
The map $\operatorname{ev}_{{\mathcal{P}}}$ in (3.2) is injective. In particular, ${\mathcal{C}}$ has dimension $k=(b-1)r$.
@@
for all $1\leq i\leq b$ and all $11\leq j\leq r+$ 1. Hence, for all $1\leq i\leq b$ the polynomials $f(x_{i},y)\in{\mathbb{F}}_{q}[y]$ of degree $\leq r-1$ have the $r+1$ distinct zeros $y_{i,1},\dots,y_{i,r+1}$, implying $f(x_{i},y)=0$ as an element of ${\mathbb{F}}_{q}[y]$. In turn, this shows that for all $10\leq\ell\leq r-$ 1 the polynomials $a_{\ell}(x)\in{\mathbb{F}}_{q}[x]$ of degree $\leq b-2$ have the $b$ distinct zeros $x_{1},\dots,x_{b}$, so $a_{\ell}(x)=0$ as an element of ${\mathbb{F}}_{q}[x]$. Hence $f(x,y)$ is the zero polynomial, as claimed, and
@@
Let $f(x,y)=\sum_{\ell=0}^{r-1}a_{\ell}(x)y^{\ell}$ be an element of $V$, where $a_{\ell}(x)\in{\mathbb{F}}_{q}[x]$ for all $10\leq\ell\leq r-$ 1. Suppose that a code word $\mathbf{c}=(f(P_{1}),\dots,f(P_{n}))$ is missing a symbol $f(P_{i})$. Without loss of generality, we may assume that $i=1$, so $P_{1}=(x_{1},y_{1,1})\in A_{1}$, where $A_{1}=\{(x_{1},y_{1,j})\}_{1\leq j\leq r+1}\subset\mathcal{P}$. Consider the matrices
@@
The code length of ${\mathcal{C}}$ is $n=b(r+1)$ by construction. By Lemma 3.3, ${\mathcal{C}}$ has dimension $k=(b-1)r$. The Singleton-type bound (2.1) gives the following upper bound for the minimum distance of ${\mathcal{C}}$
@@
and set $\mathbf{c}\coloneqq\operatorname{ev}_{\mathcal{P}}(f(x,y))$. The batches $A_{1},\dots,A_{b-2}$ are zero-fibers for $f(x,y)$, so $f(x_{i},y)\equiv 0$ in ${\mathbb{F}}_{q}[y]$ for all $21\leq i\leq b-$ 2. We also have $f(x,y_{b-1,j})\equiv 0$ in ${\mathbb{F}}_{q}[x]$ for all $11\leq j\leq r-$ 1. Moreover
@@
for $11\leq j\leq r+$ 1 because the $x$-coordinates are distinct, and the $y$-coordinates in the batch $A_{b-1}$ are distinct. Therefore $f(P)=0$ for exacly $(b-2)(r+1)+(r-1)$ points $P\in{\mathcal{P}}$. The number of nonzero entries of $\mathbf{c}$ is then
$n-((b-2)(r+1)+(r-1))=b(r+1)-(b-2)(r+1)-(r-1)=r+3.$
@@ 3.4. Minimum distance of ${\mathcal{C}}$ for small locality
If the set $A_{i}\subset{\mathcal{P}}$ is a zero-fiber for a polynomial $f(x,y)\in V$, then $f(x_{i},y)\in{\mathbb{F}}_{q}[y]$ is a polynomial of degree $\leq r-1$ with $r+1$ zeros, whence $f(x_{i},y)\equiv 0$ in ${\mathbb{F}}_{q}[y]$. Consequently $f(x_{i},y_{v,w})=0$ for all $1\leq v\leq b$ and all $11\leq w\leq r+$ 1.
@@
Let $f(x,y)=\sum_{\ell=0}^{r-1}a_{\ell}(x)y^{\ell}$ be an element of $V$, where $a_{\ell}(x)\in{\mathbb{F}}_{q}[x]$ and $\deg a_{\ell}(x)\leq b-2$ for all $10\leq\ell\leq r-$ 1. Since $f(x,y)\not\equiv 0$ in ${\mathbb{F}}_{q}[x,y]$, the code word $\mathbf{c}\coloneqq\mathrm{ev}_{\mathcal{P}}(f(x,y))$ is not zero by Lemma 3.3. Without loss of generality, we may assume $f(P_{n})\neq 0$ where $P_{n}=(x_{b},y_{b,r+1})$. Then
@@
We now prove the claim by contradiction. Assume there exist $b-1$ values $i\in\{1,\dots,b\}$ such that $f(P)=0$ for all $P\in A_{i}$. Since $f(x_{b},y_{b,r+1})\neq 0$ it must be that $f(x_{i},y_{i,j})=0$ for all $11\leq i\leq b-$ 1 and all $11\leq j\leq r+$ 1. By Observation 3.8 we have $f(x_{i},y_{v,w})=0$ for all $11\leq i\leq b-$ 1, all $1\leq v\leq b$, and all $11\leq w\leq r+$ 1. For $11\leq i\leq b-$ 1 consider
@@
These are polynomials of degree $\leq r-1$ with at least #$\#\{y_{v,w}\}_{1\leq v\leq b}^{1\leq w\leq r+1}=b(r+1)$ zeros. Since $b(r+1)>r-1$ then $f(x_{i},y)=0$ as an element of ${\mathbb{F}}_{q}[y]$ for all $11\leq i\leq b-$ 1. Thus the coefficients $a_{\ell}(x_{i})=0$ for all $11\leq i\leq b-$ 1 and all $10\leq\ell\leq r-$ 1. In particular $a_{m}(x)$ has $b-1$ zeros. This contradicts that $a_{m}(x)$ has $\leq b-2$ zeros, finishing the proof. ∎
@@
Let $b\geq 3$, let $f(x,y)\in V$ be a nonzero polynomial, and let $s\in{\mathbb{Z}}_{\geq 0}$ be number of zero-fibers of $f(x,y)$.
@@
    If $s=b-2$ then $f(P)=0$ for $\leq(b-2)(r+1)+(r-1)$ points $P\in\mathcal{P}$.
@@
By Lemmas 3.3 and 3.9, since $f(x,y)$ is not zero, we have $20\leq s\leq b-$ 2. Fix $i\in\{1,\dots,b\}$. If there exists a point $P\in A_{i}$ such that $f(P)\neq 0$, since $f(x_{i},y)\in{\mathbb{F}}_{q}[y]$ has degree $\leq r-1$, then $f(x,y)$ can vanish on $\leq r-1$ points in $A_{i}$. Thus the number of points in $\mathcal{P}$ where $f(x,y)$ vanishes is at most
$s(r+1)+(b-s)(r-1)=b(r-1)+2s,$
which increases as $s$ increases. If $s=b-2$, then, without loss of generality $f(x,y)$ vanishes at $A_{1}\amalg\cdots\amalg A_{b-2}$. In particular, for $21\leq i\leq b-$ 2, the single-variable polynomial
@@
of degree $\leq r-1$ vanishes at $y_{i,1},\dots,y_{i,r+1}$, which implies that $a_{\ell}(x)$ vanishes at $x_{1},\dots,x_{b-2}$ for all $11\leq j\leq r+$ 1, whence
@@
$(b-2)(r+1)+(r-1).$
The last part of the claim follows by observing that when $r=1,2,3$ then for all $30\leq s\leq b-$ 3 we have
@@
Note that, when $s=b-2$, we have $(b-2)(r+1)+(r-1)\leq b(r-1)+2s$.
@@
We already know from the Fiber Vanishing Lemma that $s\leq b-2$ for any nonzero code word $\mathbf{c}$. Suppose that $s=b-2$. By Lemma 3.10, $\mathbf{c}$ has weight at least
$n-((b-2)(r+1)+(r-1))=r+3.$

### v3

@@ Abstract.
A code is locally recoverable when each symbol in one of its code words can be reconstructed as a function of $r$ other symbols. We use bundles of projective spaces over a line to construct locally recoverable codes with availability; that is, evaluation codes where each code word symbol can be reconstructed from several disjoint sets of other symbols. The simplest case, where the code’s underlying variety is a plane, exhibits noteworthy properties: When $r=1$, 22, 33, they are optimal; when $r\geq 4$, they are optimal with probability approaching $1$ as the alphabet size grows. Additionally, their information rate is close to the theoretical limit. In higher dimensions, our codes form a family of asymptotically good codes.
@@ 1.2. Results
Let $r=1$, 22, or $3$ and let $b$ satisfy $3\leq b\leq\frac{q}{r+1}$. The code $\mathcal{C}$ is optimal and locally recoverable with locality $r$. Its parameters $n,k,d]_{q}$ are
@@
Let ${\mathbb{F}}_{q}$ be a finite field of cardinality $q$. Fix positive integers $\alpha$, $\beta$, $b$, and $t$ such that $b\geq\alpha+1$. There exists an integer $m_{0}=m_{0}(\alpha,\beta,b,t,q)$ such that, for each $m\geq m_{0}$, there is a locally recoverable code over ${\mathbb{F}}_{q}$ with locality $r=\binom{\beta+m}{\beta}$ and availability $t$, and parameters $n,k,d]_{q}$ satisfying
@@
The cardinality of a set $S$ will be denoted by #S\#S. A code will be denoted by a calligraphic letter $\mathcal{C}$}, and its dual code will be denoted by $\mathcal{C}$$^{\bot}$. All the codes we handle are linear evaluation codes ${\mathcal{C}}\coloneqq\mathrm{im}\operatorname{ev}_{{\mathcal{P}}}$ with $\mathcal{P}$} a finite set of points. A code word of $\mathcal{C}$} will be denoted by a bold lowercase letter $\mathbf{c}$. Scalars in ${\mathbb{F}}_{q}$ will be denoted by a lowercase $x$, and elements of ${\mathbb{F}}_{q}^{n}$ for $n>1$ will be denoted by a bold lowercase $\mathbf{x}$. The $n$-dimensional affine space over ${\mathbb{F}}_{q}$ is denoted by ${\mathbb{A}}^{n}$, and its Cartesian coordinates are denoted by ${\mathbb{A}}_{x_{1},\dots,x_{n}}^{n}$. The $n$-dimensional projective space over ${\mathbb{F}}_{q}$ is denoted by ${\mathbb{P}}^{n}$, its homogeneous coordinates are denoted by ${\mathbb{P}}_{\mathbf{x}}^{n}$ where $\mathbf{x}=[x_{0},\dots,x_{n}]$. When we write ${\mathbb{P}}_{x_{0},\dots,x_{n-1}}^{n}$, the coordinates $x_{0},\dots,x_{n-1}$ are the affine coordinates in the chart $D_{+}(x_{n})$ where the last homogeneous coordinate $x_{n}$ is nonzero. Given a projective variety $X$ over ${\mathbb{F}}_{q}$, we denote by $X({\mathbb{F}}_{q})$ the ${\mathbb{F}}_{q}$-rational points of $X$.
@@ 2.1. Linear locally recoverable codes
A code $\mathcal{C}$} is said to be locally recoverable (LRC) with locality $r$ if for each symbol $c_{i}$ in a code word $\mathbf{c}=(c_{1},\dots,c_{n})\in{\mathcal{C}}$, there exists a recovery set $R_{i}\subset\{1,\dots,n\}\setminus\{i\}$ with $R_{i}\leq r$ such that $c_{i}$ is a function of the symbols $\{c_{j}\}_{j\in R_{i}}$. In particular, if the $i$-th coordinate of $\mathbf{c}$ is lost, it can be recovered by accessing $\leq r$ other coordinates in $\mathbf{c}$. Trivially, every linear code is a LRC with locality $r=k$. Our convention of locality throughout this paper is also referred to as all-symbol locality in the literature [BargTamoVluadact17]. An LRC $\mathcal{C}$} with locality $r$ is said to have availability $t$ if for every $c_{i}$ in a code word $\mathbf{c}=(c_{1},\dots,c_{n})\in{\mathcal{C}}$ there exist $t$ disjoint recovery sets $R_{i,\ell}\subset\{1,\dots,n\}\setminus\{i\}$ with $R_{i,\ell}\leq r$ such that $c_{i}$ is a function of the symbols $\{c_{j}\}_{j\in R_{i,\ell}}$ for $1\leq\ell\leq t$.
The parameters of an LRC $\mathcal{C}$} are denoted $n,k,d;r,t]_{q}$, or simply $n,k,d;r]_{q}$ when $t=1$. They are constrained by relations like the Singleton-type bound for the minimum distance
@@
An LRC $\mathcal{C}$} with parameters $n,k,d;r]_{q}$ whose minimum distance achieves equality in (2.1) is said to be optimal. When an LRC $\mathcal{C}$} with parameters $n,k,d;r]_{q}$ has the property that each code word is partitioned into sets of $r+1$ elements where recoverability takes place within each set, as is often the case with algebro-geometric locally recoverable codes, a particularly simple proof of (2.1) was given in [SalgadoVarillyAlvaradoVoloch21]*Theorem I.3.
Let $\mathscr{C}=\{{\mathcal{C}}_{i}\}_{i=1}^{\infty}$ be a family of codes (not necessarily LRCs) and denote by $n_{i},k_{i},d_{i}]_{q}$ the parameters of ${\mathcal{C}}_{i}$ for all $i\geq 1$. We say that $\mathscr{C}$ is a family of asymptotically good codes when
@@ 2.2. Algebro-geometric evaluation codes
Algebraic Geometry supplies an abundance of constructions for linear codes under the framework of evaluation codes. To specify such a code one needs a triple $(X,{\mathcal{P}},V)$, where $X$ is a quasi-projective variety over a finite field ${\mathbb{F}}_{q}$, ${\mathcal{P}}=\{P_{1},\dots,P_{n}\}$ is a set of $n$ points in $X({\mathbb{F}}_{q})$, and $V$ is a finite-dimensional subspace of the function field of $X$. From this data, we construct a linear code as the image of the evaluation map
[…]
@@ 3.2. Relations to existing codes
for a nonnegative integer $N$. Our codes consider mostly the case where N=b2N=b-2, though we look at smaller values of $N$ in Section 3.7. The key difference between their codes and ours is the choice of points $\mathcal{P}$} for code evaluation. We all consider a set of points $\mathcal{P}$} partitioned into $b$ batches $A_{1},\dots,A_{b}$ consisting of $r+1$ distinct points. However, in [SalgadoVarillyAlvaradoVoloch21]*Section III.B each batch is additionally constrained to satisfy an extra algebraic relation: If $A_{i}=\{(x_{i},y_{i,j})\}_{j=1}^{r+1}$, then there exists $g(x)\in{\mathbb{F}}_{q}[x]$ a polynomial of degree $r+1$ such that $y_{i,j}=g(x_{i})$ for all $1\leq i\leq b$ and $1\leq j\leq r+1$. Tamo and Barg’s construction in [TamoBarg14] uses the polynomial $g(x)=x^{r+1}$, and more generally discusses the concept of good polynomials $g(x)$. The polynomial $g(x)$ affords better control of the minimum distance of the resulting codes. Without this kind of control, we can only prove that the codes constructed in Section 3.1 are optimal when $b\geq 3$ and $r=1$, 22 or $3$; see Theorem 3.13. However, when $r\geq 4$, our codes in Section 3.1 are provably optimal with probability approaching $1$ for uniformly random choices of points $\mathcal{P}$} as $q\to\infty$; see Theorem 3.15. Although our codes for $r\geq 4$ are not always optimal (see Section 4), our investigation suggests that the use of good polynomials, like $g(x)=x^{r+1}$, imposes a serious constraint on the universe’s supply of optimal LRCs.
Our codes also share superficial similarities with those of Munuera and Tenório [MunueraTenorio18]*Section 2.2, which is not surprising, as their codes generalize those in [TamoBarg14, BargTamoVluadact17]. However, our codes are neither a special case of the Munuera–Tenório construction, nor is there a clear common refinement of both constructions. To mimic our codes in the notation of [MunueraTenorio18], one would need to take $m=2$, $t=1$, $\phi_{1}=x_{1}\eqqcolon x$, $\phi_{2}=x_{2}\eqqcolon y$, ${\mathcal{S}}=b$, and $V_{i}=\{a(x)\in{\mathbb{F}}_{q}[x]:\deg a(x)\leq b-2\}$ for all $0\leq i\leq r-1$. This would force r=q1r=q-1 in their set-up (a restriction we do not impose), and the set of evaluation points ${\mathcal{P}}\subset{\mathbb{A}}^{2}_{x,y}({\mathbb{F}}_{q})$ would contain $b(r+1)=bq$ points in $b$ batches with $r+1=q$ overlapping $y$ coordinates. In our construction, it is essential that all the $b(r+1)$ points have distinct $y$ coordinates; see Remarks 3.1(3) and Figure 1.
@@ 3.3. Dimension and locality of $\mathcal{C}$}
The map $\operatorname{ev}_{{\mathcal{P}}}$ in (3.2) is injective. In particular, $\mathcal{C}$} has dimension k=(b1)rk=(b-1)r.
@@
for all $1\leq i\leq b$ and all $1\leq j\leq r+1$. Hence, for all $1\leq i\leq b$ the polynomials $f(x_{i},y)\in{\mathbb{F}}_{q}[y]$ of degree $\leq r-1$ have the $r+1$ distinct zeros $y_{i,1},\dots,y_{i,r+1}$, implying $f(x_{i},y)=0$ as an element of ${\mathbb{F}}_{q}[y]$. In turn, this shows that for all $0\leq\ell\leq r-1$ the polynomials $a_{\ell}(x)\in{\mathbb{F}}_{q}[x]$ of degree $\leq b-2$ have the $b$ distinct zeros $x_{1},\dots,x_{b}$, so $a_{\ell}(x)=0$ as an element of ${\mathbb{F}}_{q}[x]$. Hence $f(x,y)$ is the zero polynomial, as claimed, and
@@
Let $f(x,y)=\sum_{\ell=0}^{r-1}a_{\ell}(x)y^{\ell}$ be an element of $V$, where $a_{\ell}(x)\in{\mathbb{F}}_{q}[x]$ for all $0\leq\ell\leq r-1$. Suppose that a code word $\mathbf{c}=(f(P_{1}),\dots,f(P_{n}))$ is missing a symbol $f(P_{i})$. Without loss of generality, we may assume that $i=1$, so $P_{1}=(x_{1},y_{1,1})\in A_{1}$, where $A_{1}=\{(x_{1},y_{1,j})\}_{1\leq j\leq r+1}\subset\mathcal{P}$. Consider the matrices
@@
The code length of $\mathcal{C}$} is $n=b(r+1)$ by construction. By Lemma 3.3, $\mathcal{C}$} has dimension k=(b1)rk=(b-1)r. The Singleton-type bound (2.1) gives the following upper bound for the minimum distance of $\mathcal{C}$}
@@
and set $\mathbf{c}\coloneqq\operatorname{ev}_{\mathcal{P}}(f(x,y))$. The batches $A_{1},\dots,A_{b-2}$ are zero-fibers for $f(x,y)$, so $f(x_{i},y)\equiv 0$ in ${\mathbb{F}}_{q}[y]$ for all $1\leq i\leq b-2$. We also have $f(x,y_{b-1,j})\equiv 0$ in ${\mathbb{F}}_{q}[x]$ for all $1\leq j\leq r-1$. Moreover
@@
for $1\leq j\leq r+1$ because the $x$-coordinates are distinct, and the $y$-coordinates in the batch $A_{b-1}$ are distinct. Therefore $f(P)=0$ for exacly (b2)(r+1)+(r1)(b-2)(r+1)+(r-1) points $P\in{\mathcal{P}}$. The number of nonzero entries of $\mathbf{c}$ is then
n((b2)(r+1)+(r1))=b(r+1)(b2)(r+1)(r1)=r+3.n-((b-2)(r+1)+(r-1))=b(r+1)-(b-2)(r+1)-(r-1)=r+3.
@@ 3.4. Minimum distance of $\mathcal{C}$} for small locality
If the set $A_{i}\subset{\mathcal{P}}$ is a zero-fiber for a polynomial $f(x,y)\in V$, then $f(x_{i},y)\in{\mathbb{F}}_{q}[y]$ is a polynomial of degree $\leq r-1$ with $r+1$ zeros, whence $f(x_{i},y)\equiv 0$ in ${\mathbb{F}}_{q}[y]$. Consequently $f(x_{i},y_{v,w})=0$ for all $1\leq v\leq b$ and all $1\leq w\leq r+1$.
@@
Let $f(x,y)=\sum_{\ell=0}^{r-1}a_{\ell}(x)y^{\ell}$ be an element of $V$, where $a_{\ell}(x)\in{\mathbb{F}}_{q}[x]$ and $\deg a_{\ell}(x)\leq b-2$ for all $0\leq\ell\leq r-1$. Since $f(x,y)\not\equiv 0$ in ${\mathbb{F}}_{q}[x,y]$, the code word $\mathbf{c}\coloneqq\mathrm{ev}_{\mathcal{P}}(f(x,y))$ is not zero by Lemma 3.3. Without loss of generality, we may assume $f(P_{n})\neq 0$ where $P_{n}=(x_{b},y_{b,r+1})$. Then
@@
We now prove the claim by contradiction. Assume there exist b1b-1 values $i\in\{1,\dots,b\}$ such that $f(P)=0$ for all $P\in A_{i}$. Since $f(x_{b},y_{b,r+1})\neq 0$ it must be that $f(x_{i},y_{i,j})=0$ for all $1\leq i\leq b-1$ and all $1\leq j\leq r+1$. By Observation 3.8 we have $f(x_{i},y_{v,w})=0$ for all $1\leq i\leq b-1$, all $1\leq v\leq b$, and all $1\leq w\leq r+1$. For $1\leq i\leq b-1$ consider
@@
These are polynomials of degree $\leq r-1$ with at least $\{y_{v,w}\}_{1\leq v\leq b}^{1\leq w\leq r+1}=b(r+1)$ zeros. Since b(r+1)>r1b(r+1)>r-1 then $f(x_{i},y)=0$ as an element of ${\mathbb{F}}_{q}[y]$ for all $1\leq i\leq b-1$. Thus the coefficients $a_{\ell}(x_{i})=0$ for all $1\leq i\leq b-1$ and all $0\leq\ell\leq r-1$. In particular $a_{m}(x)$ has b1b-1 zeros. This contradicts that $a_{m}(x)$ has $\leq b-2$ zeros, finishing the proof. ∎
@@
Let $b\geq 3$, let $f(x,y)\in V$ be a nonzero polynomial, and let $s\in{\mathbb{Z}}_{\geq 0}$ be the number of zero-fibers of $f(x,y)$.
@@
    If s=b2s=b-2 then $f(P)=0$ for $\leq(b-2)(r+1)+(r-1)$ points $P\in\mathcal{P}$.
@@
By Lemmas 3.3 and 3.9, since $f(x,y)$ is not zero, we have $0\leq s\leq b-2$. Fix $i\in\{1,\dots,b\}$. If there exists a point $P\in A_{i}$ such that $f(P)\neq 0$, since $f(x_{i},y)\in{\mathbb{F}}_{q}[y]$ has degree $\leq r-1$, then $f(x,y)$ can vanish on $\leq r-1$ points in $A_{i}$. Thus the number of points in $\mathcal{P}$ where $f(x,y)$ vanishes is at most
s(r+1)+(bs)(r1)=b(r1)+2s,s(r+1)+(b-s)(r-1)=b(r-1)+2s,
which increases as $s$ increases. If s=b2s=b-2, then, without loss of generality $f(x,y)$ vanishes at $A_{1}\amalg\cdots\amalg A_{b-2}$. In particular, for $1\leq i\leq b-2$, the single-variable polynomial
@@
of degree $\leq r-1$ vanishes at $y_{i,1},\dots,y_{i,r+1}$, which implies that $a_{\ell}(x)$ vanishes at $x_{1},\dots,x_{b-2}$ for all $1\leq j\leq r+1$, whence
@@
(b2)(r+1)+(r1).(b-2)(r+1)+(r-1).
The last part of the claim follows by observing that when $r=1,2,3$ then for all $0\leq s\leq b-3$ we have
@@
Note that, when s=b2s=b-2, we have $(b-2)(r+1)+(r-1)\leq b(r-1)+2s$.
@@
We already know from the Fiber Vanishing Lemma that $s\leq b-2$ for any nonzero code word $\mathbf{c}$. Suppose that s=b2s=b-2. By Lemma 3.10, $\mathbf{c}$ has weight at least
n((b2)(r+1)+(r1))=r+3.n-((b-2)(r+1)+(r-1))=r+3.